Sickle-cell anaemia is an example of how fitness depends on environmental conditions. In places where malaria, caused by Plasmodium falciparum is common, HbAHbS heterozygotes have a selective advantage compared to individuals who are homozygotes for the wild type allele. However, HbSHbS homozygotes suffer from sickle-cell anaemia and usually die before producing offspring. The relative fitness (w) of the three genotypes in Nigeria are:

HbAHbA 0.78,

HbAHbS 1.0

HbSHbS 0.07

a. What is the mathematical relationship between relative fitness and selection coefficient?

b. What is the selection coefficient (s) for each genotype?

c. Calculate the frequency of the three genotypes after one generation of selection, if the genotype frequencies before selection (i.e. at birth) are 0.4 for HbAHbA homozygotes, 0.46 for heterozygotes and 0.14 for HbSHbS homozygotes.

d. Is it likely that over the next 100 years the HbS allele will be lost in the population?
Answer #1

a. Relation between relative fitness and selection coefficient-

s = 1 - w.

b.

HbAHbA 1 - 0.78 = 0.22,

HbAHbS 1 - 1 = 0

HbSHbS 1 - 0.07 = 0.93

c. For HbA-

p = 1 / 0.22 + 1 = 0.82

For HbS-

q = 1 - 0.82 = 0.18

d. The allele HbS will not get lost, as it will be masked in the genotype HbA HbS, which is the most favored genotype and helps in survival of population.